OpenStudy (rocioballonxox):

What value of x makes this sentence true? 5x + 12 = x – 8

OpenStudy (.mehek.):

Move x to one side and 12 to the other side 5x - x = ? -8 - 12 = ?

OpenStudy (plainntall):

Remember you can do what Mehek says by doing the same thing to both sides of the equation 5x+12=x-8 5x+12-12 = x-8-12 (subtract 12 from both sides 5x+0=x-20 5x-x=x-x-20 (substract x from both sides) 4x=-20 4x/4=-20/4 (divide both sides by 4) x=-5 checking we have 5(-5)+12=(-5)-8 -25+12=-13 -13=-13 it checks out